OpenStudy (anonymous):

V=((1/3)(pi)(r^2)(h))(3r-h) solve for h

jhonyy9 (jhonyy9):

you have thought it in this way ? pi*r^2 *h*(3r-h) V = ----------------- 3 - so than this is right will get for h : 3V = pi*r^2*h*(3r -h) 3V = 3pi*r^3*h - pi*r^2*h^2 pi*r^2*h^2 -3pi*r^3*h +3V = 0 - so in this case than you see you have got a quadratic equation in ,,h" - what you can solving exactly like a quadratic equation with unknown x using discriminant - hope this will help you